From Darin Koch:
Friday, December 10.
82 guests total. Opened observatory at 6 p.m., closed at 10:30 p.m.
Both telescopes working well. Grass mowed. Computer still not working.
Thursday, December 9.
Eight guests total. Opened observatory at 5:50 p.m., closed at 10:15 p.m.
A 40mm eyepiece was not put away, left in Harlan Smith cabinet with O III filter still attached. Ealing Telrad was not turned off. Collimated the Harlan Smith telescope.
Both telescopes working well, Internet working, grass mowed.